Status Report

NASA ISS On-Orbit Status 20 June 2009

By SpaceRef Editor
June 20, 2009
Filed under , , ,
NASA ISS On-Orbit Status 20 June 2009

All ISS systems continue to function nominally, except those noted previously or below. Saturday – light-duty day for CDR Gennady Padalka (Russia), FE-1 Michael Barratt (USA), FE-2 Koichi Wakata (Japan), FE-3 Roman Romanenko (Russia), FE-4 Robert Thirsk (Canada), FE-5 Frank DeWinne (Belgium).

Thirsk & DeWinne began the last day continuing their second session with the NASA/JSC experiment NUTRITION w/Repository, collecting urine samples throughout the day for storage in the MELFI (Minus-Eighty Laboratory Freezer for ISS). [The NUTRITION project is the most comprehensive in-flight study done by NASA to date of human physiologic changes during long-duration space flight. It includes measures of bone metabolism, oxidative damage, nutritional assessments, and hormonal changes, expanding the previous Clinical Nutritional Assessment profile (MR016L) testing in three ways: Addition of in-flight blood & urine collection (made possible by supercold MELFI dewars), normative markers of nutritional assessment, and a return session plus 30-day (R+30) session to allow evaluation of post-flight nutrition and implications for rehabilitation.]

The six-member crew performed the regular weekly three-hour task of thorough station cleaning, including COL (Columbus Orbital Laboratory) and Kibo. ["Uborka", usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the SM (Service Module) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the sleep stations with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.]

As part of the house cleaning, Padalka & Romanenko conducted regular maintenance inspection & cleaning on fan screens, Group A, in the FGB (TsV2), DC1 (V3), and SM (VPkhO, VPrK, FS5, FS6 & FS9), plus dust filter replacement in the FGB.

Later, CDR Padalka performed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM, including the weekly collection of the toilet flush (SP) counter and water supply (SVO) readings for calldown to TsUP-Moscow.

FE-4 Thirsk filled out his third regular weekly FFQ (Food Frequency Questionnaire) on the MEC (Medical Equipment Computer). [On the FFQs, NASA astronauts keep a personalized log of their nutritional intake over time on special MEC software. Recorded are the amounts consumed during the past week of such food items as beverages, cereals, grains, eggs, breads, snacks, sweets, fruit, beans, soup, vegetables, dairy, fish, meat, chicken, sauces & spreads, and vitamins. The FFQ is performed once a week to estimate nutrient intake from the previous week and to give recommendations to ground specialists that help maintain optimal crew health. Weekly estimation has been verified to be reliable enough that nutrients do not need to be tracked daily.]

At ~9:05am EDT, the crew conducted their regular WPC (Weekly Planning Conference) with the ground, discussing next week’s "Look-Ahead Plan" (prepared jointly by MCC-Houston and TsUP-Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.

Koichi Wakata performed the regular bi-monthly reboots of the OCA Router and File Server SSC (Station Support Computer) laptops.

The crew completed their regular daily 2.5-hr. physical workout program on the CEVIS cycle ergometer (FE-1, FE-2), TVIS treadmill with vibration isolation (CDR, FE-3, FE-4, FE-5), ARED advanced resistive exercise device (CDR, FE-1, FE-2, FE-4, FE-5) and VELO cycle ergometer with bungee cord load trainer (FE-3). [On the CEVIS, the actual loads remain slightly lower than the commanded loads, but this was expected. A manual correction of the pertinent calibration coefficient via the control panel touch screen will be done at a later time when the new value has been determined.]

Later, Mike Barratt transferred the exercise data file to the MEC (Medical Equipment Computer) for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on ARED, followed by their erasure on the HRM storage medium (done six times a week).

Weekly Science Update (Expedition Twenty — Week 3)

3-D SPACE: “Frank, good work on your first 3D-SPACE session on 6/17! Bob, good work on your first 3D-SPACE session on 6/18! We look forward to receiving the crew note you left on 3D-SPACE.”

AgCam (Agricultural Camera): “Mike, we are analyzing the data to figure out what’s wrong with our system. We appreciate all your help in the past, and look forward to your support in the future. We are currently putting together a troubleshooting procedure and we are hoping get it ready for you soon, potentially as early as next week.

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Complete.

ALTEA DOSI (NASA/ASI): Standing by.

BCAT-4/5 (Binary Colloidal Alloy Test 4/5): BCAT-4 ongoing. BCAT-5 to be launched on 2J/A.

BIOLAB (ESA): No report.

Biological Rhythms (JAXA): “Bob. your first measurement for Bio-Rhythms was completed and we have received the downlinked data. Your second measurement is planned for August. Frank, your first measurement was started this week on GMT 169. Your second measurement will be planned for July.”

BISE (CSA, Bodies in the Space Environment): “Bob and Frank, we would like to first thank you for completing another successful run. The data sets look very good and are we are very happy with the exciting run that we had this week.”

BISPHOSPHONATES: “Koichi, thanks for completing your pill ingestion. Your next session is scheduled for 6/22. Bob, thanks for completing your pill ingestion. Your next session is also scheduled for 6/22.”

CARD (Long Term Microgravity: A Model for Investigating Mechanisms of Heart Disease, ESA): “Thank you, Koichi, for performing this overnight BP (blood pressure) measurement and also for doing the parallel measurement with the BP-ECG. The overnight measurement will allow for the completion of your data from your CARD session.”

CARDIOCOG-2: Complete.

CBEF (JAXA Cell Biology Experiment Facility): Regular maintenance task was completed this week.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): “Mike, initial analysis of your second CCISS session is complete. The data looks good and we will be sending you a private message with the preliminary results. Thanks again for repeating the Baro and 24-HR Heart Rate Study. Bob, congratulations on completing your first on-orbit CCISS session! Thanks for repeating the Baro Study and catching the full Holter card. We were hoping it would last until the completion of the CCISS activity. Dr. Hughson supported console from Waterloo and looks forward to getting the data.”

CFE (Capillary Flow Experiment): Reserve.

CSI-3/CGBA-5 (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus 5): Complete.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CIR (Combustion Integrated Rack): MDCA/Flex: “This week we plan to rerun one previous heptane test point and run one new heptane test point. We certainly appreciate the effort to move the “water wall” away from contact with the CIR so we can determine what our true microgravity environment looks like and see if our earlier test points might have been influenced by the objects being against the rack. If the two test points are successful this time, we will be able to obtain a better idea of what the microgravity influences are on droplet deployment. Thanks again!”

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

Commercial 2 (JAXA): Completed.

Commercial 3 (JAXA): Completed.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Complete.

DomeGene (JAXA): Complete.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Planned.

EDR (European Drawer Rack, ESA): The rack is continuously active in support of the PCDF (Protein Crystallization Diagnostic Facility) experiment. EDR is providing power / data and temperature control (via cooling loop) to PCDF.

ELITE-S2 (Elaboratore Immagini Televisive – Space 2): Planned.

ENose (Electronic Nose): Operating.

EPM (European Physiology Module): Activated on 6/18 to support CARD activities.

EPO (Educational Payload Operations): Reserve.

EPO 3-min Video (JAXA): Self video shooting for 10 science missions completed by Koichi. MPC downlink was also completed.”

EPO J-Astro Report (JAXA): Ongoing.

EPO Space Clothes (JAXA): Complete.

EPO Hiten (Dance, JAXA): Complete.

EPO Moon Score (JAXA): Planned.

EPO Try Zero-G (JAXA): “VS #3 (Two-some: arm-wrestling, a handshake, a pushing technique and a tug of war) has been completed on last weekend. With the cooperation of Robert Thirsk from CSA, Koichi showed great performances to demonstrate the principle of action and reaction in 0G in addition to his solo performances. Their witty performances made the media people laugh and we really appreciate Koichi and Robert’s skillful camera arrangement. Thanks to their efforts, this educational and media opportunity was a great success.”

EPO Kibo Kids Tour (JAXA): Complete.

EPO Spiral Top (JAXA): “Additional photo was very interesting. Thank you. Please keep photo in proprietary folder when you download them.”

ETD (Eye Tracking Device): Completed.

EuTEF (European Technology Exposure Facility): EuTEF platform and its instruments have been running nominally during this reporting period.– DEBIE-2: The instrument continues to regularly generate empty science packets, and as a work-around the instrument script includes a daily power cycle command, which is skipped from ground if the science packets remain nominal. A new script was started on 6/17. — DOSTEL, EXPOSE, MEDET: On-going science acquisition – nominal;- EVC: Unsuccessful attempts to operate EVC due to High Rate Data and over-current problems;– FIPEX: A new EOP script was started on 6/17 and will continue running till 6/21; — PLEGPAY: Instrument is currently on but not acquiring science; — TRIBOLAB: On 6/12, an additional test of the Ball Bearing 1 experiment was carried out, unfortunately it was not successful.

FACET (JAXA): “Experiment number from 35 to 39 has finished. All experiments were complete yesterday (6/12) as planned. Nucleation and homogenization are going well.”

FSL (Fluid Science Laboratory): 24hrs MMA measurements have been carried out between 6/17 and 6/18 to support PCDF EP2 Cycle 3 run.

GEOFLOW: No report.

HDTV System (JAXA): To be launched by HTV1.

Holter ECG (JAXA): Complete.

HQPC (JAXA): To be launched by 34P.


ICV (Integrated Cardiovascular): “Bob, the Integrated Cardiovascular team is very excited to be rounding out the first ICV in-flight session! Thanks for going above and beyond in your efforts to ensure a flawless data collection session. We really appreciate the insight you’ve been providing us throughout this activity. From the Holter electrode placement video (wonderful modelling job, by the way!) to your OSTPV crew notes (we are avid readers!), the extra information is very helpful to us! Thanks also for your concern about our science and hardware (double-checking the HM2 display during event marking, requesting CEVIS vs.TVIS for exercise, etc.). Awesome job!”

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): Complete.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Completed.

Marangoni Experiment for ISS in JAXA FPEF (Fluid Physics Experiment Facility): In progress.

MAXI (JAXA): Ongoing.

MDCA/Flex: “We experienced unexpected surface tension issues with Heptane fuel at 3 atmosphere pressure that preclude free deployment of the fuel. These conditions will need to be conducted with a support fiber. Test point execution order is being evaluated to best accommodate this change.”

Micro-G Clay (JAXA EPO): Complete.

MMA (Microgravity Measurement Apparatus): “Third Micro G measurement for TCQ (Temporary Crew Quarters) was performed to investigate how microgravity disturbance will transmit to the Ryutai rack. Data is now being investigated. We caught Koichi’s tapping clearly by MMA sensor and it is good reference for analysis.”

MISSE (Materials ISS Experiment): Ongoing.

Moon Photography from ISS (JAXA EPO): One run performed on the last day of Increment 18.

MSG-SAME (Microgravity Science Glovebox): Complete.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 “Pirs”.

MULTIGEN-1: Completed.

NEUROSPAT (ESA/Study of Spatial Cognition, Novelty Processing and Sensorimotor Integration): Frank performed a data transfer between EPM and the MPPLL on 6/5. This was successfully completed and there is now a back-up of the NEUROSPAT data from both sessions. “Frank, thank you for making it possible for us to monitor the data transfer progress via video.”

NOA-1/-2 (Nitric Oxide Analyzer, ESA): Complete.


PADLES (JAXA, Area PADLES 3; Passive Dosimeter for Lifescience Experiment in Space): Continuing radiation dose accumulation.

PCDF-PU (Protein Crystallization Diagnostic Facility – Process Unit): EP2Cycle2 was stopped on 6/11, allowing science team to assess the results. Meanwhile, some additional planned runs have been carried out with EP3 reactor, i.e. EP3 Cycle 5 from 6/11 to 6/14, and EP3 Cycle 6 from 6/14 to 06/17. These runs should normally conclude the experimental objectives on this EP3 reactor. Since 6/17, a new run with EP2 reactor has been started and it will be executed automatically for several days. This run will aim at studying the nucleation processes which lead to subsequent protein crystal growth. With the delay of 2J/A (STS-127) launch, the PCDF science team is currently assessing the strategy to finalize the original science objectives and if possible to add extra runs until return with 2J/A. The EP4 reactor still contains some nice crystals to be returned by 2J/A.

PCRF (Protein Crystallization Research Facility) Reconfiguration (JAXA): Complete.

PMDIS (Perceptual Motor Deficits in Space): Complete.


RadGene & LOH (JAXA): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Complete.

SEDA-AP (JAXA): Exposed Payload,- to be launched by 2JA.

SHERE (Shear History Extensional Rheology Experiment): Complete.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): “Koichi, thanks for completing your last week of in-flight sleep logging! We will be down linking the data for the PI later this month. Your only remaining in-flight activity for the Sleep experiment is doffing your Actiwatch on the Shuttle. Mike, thanks for completing another week of sleep logging. We appreciate your participation. And thanks for downloading and initializing all the Actiwatches.”

SMILES (JAXA): Exposed Payload, to be launched by HTV1.

SOLAR (Solar Monitoring Observatory): The Sun visibility window #16 ended on 6/4. The SOLAR platform has encountered an Analog Input Board anomaly (known and recurrent one). The next Sun visibility window (#17) has started on 6/13. At the beginning of the current Sun observation window, the SOLAR platform went to Stand-By Mode (known and recurrent one). SOLAR got recovered and science acquisition is proceeding nominally. The current Sun observation window is planned to end by 6/25.—SOLSPEC: science acquisition is on-going.—SOLACES: science acquisition is on-going

SOLO (Sodium Loading in Microgravity): No report.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): Reserve.

SPICE (Smoke Point In Co-flow Experiment): “You successfully performed the soot removal activities on Monday which resulted in good measurements for our runs. Your runs look good and we obtained good images of the lifted flames. Our whole team would like to thank you for your hard work this week and we look forward to having some exciting runs next week.”

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.


VLE (Video Lessons ESA): VLE-1 completed.

WAICO #1/#2 (Waving and Coiling of Arabidopsis Roots at Different g-levels): Complete/Planned (2J/A Stage).

CEO (Crew Earth Observations): Through GMT Day 168, the ground has received a total of 13,820 of ISS CEO imagery for review and cataloguing. “ We are pleased to confirm your acquisition of views of the following targets: Madeira Island, Portugal – excellent imagery, nearly cloud-free – target complete; Coast Mountains, BC, Canada – too many clouds – we’ll try again; Baikonur Cosmodrome – 20 images are under review; Madrid, Spain – incomplete; and Istanbul, Turkey – 20 frames under review. The quality, content, and composition of your photography remain high and your keen interest in CEO continues to acquire many interesting views of our Earth. Your vigilance late last week netted exciting views of an eruption of the Sarychev Volcano in Russia’s Kuril Islands northeast of Japan on GMT day 163. One of those frames will be published on NASA/GSFC’s Earth Observatory website this weekend. It shows the early stages of an explosive, volcanic eruption plume over Matua Island with indications of pyroclastic flow descending the northeast slope of Sarychev. Outstanding dynamic event imagery!”]

CEO photography can be studied at this “Gateway” website: (as of 9/1/08, this database contained 770,668 views of the Earth from space, with 324,812 from the ISS alone).

ISS Orbit (as of this morning, 7:43am EDT [= epoch])
Mean altitude — 347.8 km
Apogee height – 353.7 km
Perigee height — 341.9 km
Period — 91.49 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0008758
Solar Beta Angle — -2.3 deg (magnitude decreasing)
Orbits per 24-hr. day — 15.74
Mean altitude loss in the last 24 hours — 73 m
Revolutions since FGB/Zarya launch (Nov. 98) — 60647

SpaceRef staff editor.